OpenStudy (anonymous):

find the limit of lim x->3 (x-3)/(absx-3)

myininaya (myininaya):

\[\lim_{x \rightarrow 3}\frac{x-3}{|x|-3}\] ?

OpenStudy (anonymous):

left hand limit = lim x->3 (x-3)/-(x-3)=-1 right hand limit=lim x->3 (x-3)/(x-3)=+1 as LHL != RHL i think limit will not exist .

OpenStudy (anonymous):

absolute value of (x-3)

myininaya (myininaya):

\[x>0 =>|x|=x\] \[\lim_{x \rightarrow 3}\frac{x-3}{|x|-3}=\lim_{x \rightarrow 3}\frac{x-3}{x-3}=1\]

myininaya (myininaya):

ok prashant did it for abs(x-3)

OpenStudy (anonymous):

right hand side limit and left side limit are equal when you approaching from left side as 2.999 it'll not be -x so in both case ans will be 1

myininaya (myininaya):

no prashant is right the limit does not exist

OpenStudy (anonymous):

if x=2.99 2.99-3/2.99-3=1 if x=3.011 3.011-3/3.011-3=1 what are you talking myiniaya it happens if x was appraoching zero

myininaya (myininaya):

x is approaching 3

OpenStudy (anonymous):

so as 2.99 it's not less than zero mode (x)=x ,x>0

myininaya (myininaya):

\[\lim_{x \rightarrow 3^-}\frac{x-3}{|x-3|}=\lim_{x \rightarrow 3} \frac{x-3}{-(x-3)}=-1 \neq 1=\lim_{x \rightarrow 3} \frac{x-3}{x-3}=\lim_{x \rightarrow 3^+}\frac{x-3}{|x-3|}\] since left limit doesn't equal right limit the limit does not exist

myininaya (myininaya):



myininaya (myininaya):

but you are doing another problem right? you said if we are approaching 0?

OpenStudy (anonymous):

it is mode (x) -3

myininaya (myininaya):

not he said abs(x-3)